### v1

# SADG: Segment Any Dynamic Gaussian Without Object Trackers

Abstract
Understanding dynamic 3D scenes is fundamental for various applications, including extended reality (XR) and autonomous driving. Effectively integrating semantic information into 3D reconstruction enables holistic representation that opens opportunities for immersive and interactive applications. We introduce SADG, Segment Any Dynamic Gaussian Without Object Trackers, a novel approach that combines dynamic Gaussian Splatting representation and semantic information without reliance on object IDs. In contrast to existing works, we do not rely on supervision based on object identities to enable consistent segmentation of dynamic 3D objects. To this end, we propose to learn semantically-aware features by leveraging masks generated from the Segment Anything Model (SAM) and utilizing our novel contrastive learning objective based on hard pixel mining. The learned Gaussian features can be effectively clustered without further post-processing. This enables fast computation for further object-level editing, such as object removal, composition, and style transfer by manipulating the Gaussians in the scene. We further extend several dynamic novel-view datasets with segmentation benchmarks to enable testing of learned feature fields from unseen viewpoints. We evaluate SADG on proposed benchmarks and demonstrate the superior performance of our approach in segmenting objects within dynamic scenes along with its effectiveness for further downstream editing tasks. Our project page is available at: https://yunjinli.github.io/project-sadg/.
1 Introduction
We live in a dynamic, three-dimensional world. Most of the time, we describe our environment at a high level of objects. As we move around and observe the scene from different viewpoints, we complete the geometry and perceive the colors contributing to our object representation. At the same time, we infer dynamic properties with object motion and deformations. Despite pose, color, and shape changes, the underlying semantic meaning stays constant. This intuition can serve as a strong constraint for neural 3D representations that have gained recent popularity and promote innovation in augmented reality, gaming, and autonomous driving interactive applications. In these technologies, seamlessly integrating semantic information into underlying representation ensures an immersive experience and gives users advanced control over the dynamic scene contents.
Neural Radiance Fields (NeRFs) have revolutionized the field of 3D reconstruction and demonstrated excellent performance in synthesizing photorealistic novel views of a scene from any angle. The limitations of vanilla NeRFs promoted many follow-up works, including those that encompass semantic information by lifting 2D segmentation masks [24, 7, 19] or distilling features [21, 18] and those that faithfully represent dynamic scenes [41, 35, 27]. Due to continuous implicit representation, these methods are computationally intensive as they require re-rendering from multiple views to ensure edit consistency, which hinders their applicability for interactive downstream tasks.
Building on the strengths of NeRFs, Gaussian splatting (3DGS) [17] has introduced a powerful new approach that prioritizes efficiency and speed. Dynamic versions of 3DGS [30, 56] have offered faithful reconstruction of scenes with motion. These models can serve as a geometric prior for learning semantic features that are consistent in both space and time and open new possibilities for real-time interactive applications like scene editing and dynamic interactions. Towards this, a line of works has extended geometric properties of colored and explicit Gaussians with semantic information [6, 57], producing consistent novel view synthesis and segmentation. Nonetheless, these works are limited to static scenes and cannot handle objects in motion. There are only a few concurrent works [16, 25], which unify semantics and dynamics to comprehensively understand real-world environments. However, SA4D [16] relies on the supervision of object trackers and suffers from semantic identity conflicts in multi-view scenes. Moreover, their method is sensitive to mask noise and requires post-processing to remove degenerate Gaussians. At the same time, DGD [25] has a higher memory requirement and longer training time as they are learning 4D feature fields based on 512-dimensional CLIP [42] or 384-dimensional DINOv2 [34]. Our work aims to tackle both single- and multi-view representations with fast rendering times, minimal post-processing, and low memory footprint.
To this end, we introduce our novel framework, Segment Any Dynamic Gaussian Without Object Trackers (SADG), which effectively combines dynamic Gaussian splatting backbone [56] and semantic information without any reliance on trackers that maintain consistent object identities. To learn the spatio-temporal semantic field, we define a new contrastive objective based on hard positive and negative mining with the Segment Anything Masks (SAM) [20]. We inherit the strengths of zero-shot 2D scene understanding of the foundation model to learn the semantically-aware feature field and produce consistent spatio-temporal segmentation. Our 32-dimensional features are compact and do not depend on features of a 2D segmentation model. Rendering time and storage requirements are almost unaffected, enabling real-time interaction and editing. The learned feature field can be effectively clustered without any further post-processing and utilized in a range of downstream editing tasks, including object style transfer [14], re-coloring, composition, and removal. Given an early stage of development and benchmarks for semantic segmentation for dynamic scenes, we propose to extend existing novel-view synthesis datasets with semantic benchmarks. The proposed evaluation protocol includes single- and multi-view scenes with challenging motions and diverse contents, allowing extensive testing of semantically-aware latent representations from unseen viewpoints.
Our contributions can be summarized as follows:
We propose Segment Any Dynamic Gaussian Without Object Trackers (SADG), a new approach that achieves multi-view consistent segmentation of dynamic scenes without any tracking supervision.
We leverage information from 2D masks to learn semantically-aware latent representation for dynamic scenes using a novel contrastive learning objective.
We extensively evaluate our method on single- and multi-view scenes across five dynamic novel-view benchmarks and demonstrate state-of-the-art segmentation performance.
We demonstrate the general nature of our feature space and apply it to several downstream tasks, including object removal, style transfer, and scene composition.
We offer an interactive framework that allows to edit the scene with a simple mouse click or a text prompt. Our tool does not depend on input IDs or 2D strokes and operates in real-time.
2 Related Works
2.1 Semantic Feature Learning
With the emergence of various vision foundation models, such as Segment Anything Model (SAM) [20], DINO [5], and DINOv2 [34], researchers start to integrate the semantics from these foundation models into their reconstruction. SA3D [7] is the pioneering method that extends 2D SAM masks across images into 3D consistent object masks. GARField [19] introduces hierarchical grouping, leveraging the physical scales of 2D masks to group objects of different sizes and represent scenes at different granularity scales. Gaussian Grouping [57] is a pioneering Gaussian-based approach that utilizes consistent object IDs generated by video tracker DEVA [8] to render object IDs into camera views. Concurrent work such as SAGA [6] utilizes the masks generated from SAM, combined with scales obtained from 3D data inspired by GARField, to segment objects of various sizes. Notably, these methods are designed for static scenes. To date, only two concurrent work, SA4D [16] and DGD [25], addresses the segmentation task in dynamic 3DGS. DGD extends the rasterization pipeline of 3DGS to be able to render semantic feature of each Gaussian. The rendered Gaussian feature map is then trained with the supervision of DINOv2 [34] or CLIP [42] to mimic the output from their feature extractors. The biggest drawback of such supervision is anticipated longer training times as the dimensions of the features generated by DINOv2 and CLIP are quite large (384 and 512). Similar to Gaussian Grouping and its derivatives [11, 31], SA4D relies on video trackers to associate masks across different views to produce consistent object IDs. These approaches face significant limitations, as the mask association often fails to offer consistent object IDs in dynamic scenes, such as the oven-mitts sequence from HyperNeRF [36] as shown in Fig. 2.
2.2 Scene Editing Applications
Learning semantically-aware latent space opens opportunities for interactive downstream applications, allowing users to manipulate scenes on an object level. Several works proposed editing techniques for radiance field representations [13, 49], which are limited to a single object. On the other hand, scene-based methods [53, 24] learn a separate radiance field per object and render the full scene by combining learned opacities along the viewing rays. While the structure is modular, its editing application is limited to a pre-defined number of objects. Our method is guided by user interaction and is capable of operating on any object in the scene. Another line of work distills features such as DINO [34] into volumetric representation and shows editing capabilities on learned continuous feature field [48]. Due to the implicit representation, scene editing requires a re-assessment of the final consistency by rendering from novel scenes, which impedes the model’s applicability to real-time interactive applications.
At the same time, Gaussian splatting methods mitigate the runtime limitation of NeRFs with their explicit point-based representation. SAGA [57] proposes Local Gaussian Editing scheme, which localize the Gaussians and manipulate them according to the editing task. Feature3DGS [60] distills features from 2D foundation models such as SAM [43] and LSeg [26] enabling promptable and language-driven 3D edits. Nonetheless, these methods operate only on static scenes, while our method is capable of editing dynamic objects and ensuring temporal consistency of the edits. DGD [25] shows semantic texture editing performance by utilizing a pre-trained diffusion model and fine-tuning dynamic 3DGS representation with the reconstruction loss. On the other hand, we demonstrate a variety of editing tasks, including style transfer, object removal, and composition. Despite offering several scene editing opportunities, SA4D [16] has limited user interactivity and requires prior knowledge of object identities. Our framework contains a user-friendly graphical interface that allows operation with a simple text prompt or a mouse click.
3 Method
In this section, we introduce our novel framework SADG. As illustrated in Fig. 1 the pipeline comprises two main components: dynamic geometry reconstruction (Sec. 3.1) and Gaussian feature learning (Sec. 3.2). We adopt the Deformable-3DGS [56] pipeline to reconstruct the 4D scene. Once the 4D reconstruction is learned, we freeze this representation and proceed with Gaussian feature learning using SAM [20] masks and our novel contrastive learning objective in the rendered feature space.
3.1 Dynamic Geometry Reconstruction
We follow the approach proposed in [56] and describe it briefly for completeness. Unlike Dynamic3DGS [30] model, where the reconstruction is stored per frame, resulting in high memory consumption, an MLP learns per-Gaussian deformation with respect to the static canonical space , which is defined as
is the center position of the i-th Gaussian, is the quaternion representing the rotation, is the scaling vector, refers to the opacity, and is the Spherical Harmonic Coefficients encoding the color information () as per [56].
The resulting Gaussians at timestamp are defined as
Once is obtained, we proceed with the standard rasterization pipeline [17] to render the image . The color loss is computed with the ground truth image as
where refers to the weighting parameter for the structural similarity loss term.
3.2 Gaussian Feature Learning
Problem statement Most of the existing works [16, 57] utilize a video tracker to provide consistent object mask IDs across views and render the segmentation field directly by predicting the per-pixel probability of different object classes. However, such approaches have a significant bottleneck, namely their reliance on perfect object mask association across views. When mask inconsistencies of the same object occur between views, the entire optimization pipeline can break down, leading to sub-optimal results or failure.
Therefore, we propose a method that aims to learn the 3D semantics of objects and leverage them directly for downstream tasks across different views. In other words, instead of predicting per-pixel probabilities during rendering, we focus on integrating semantic information into Gaussians and leverage semantic constancy for learning multi-view consistent spatio-temporal feature field.
Representation Given 4D reconstruction, we incorporate semantic information into the scene. Specifically, we extend its -th Gaussian with a 32-dimensional semantic feature vector , the so-called “Gaussian feature”. The semantically-aware feature field of a scene represented with Gaussians can be expressed as
Rendering Gaussian Features The learned Gaussian features can be rendered similarly to [17, 60] with point-based -rendering [23, 22]. Given a sorted subset of Gaussians that are related to pixel , the resulting rendered features can be computed as follows:
Existing methods [6] use smooth Gaussian features instead, as they help to filter out noisy 3D Gaussian features with high similarity scores corresponding to unrelated objects. The smooth Gaussian feature for the -th Gaussian is computed as per Eq. 6, where denotes a set of nearest neighbors [9] of the -th Gaussian based on its 3D position. We adopt this convention in our work and utilize smooth feature vectors in the rendering process (Eq. 5).
Our contrastive semantically-aware learning objective
The generated set of SAM masks comprise binary masks of a given image , where each mask corresponds to a different object. In practice, we do not consider all masks from SAM to alleviate memory and computational requirements. Instead, we randomly sample masks for further usage.
We define the pixel-mask correspondence vector in Eq. 7, which captures mask information for a certain pixel coordinate (, ) across all binary masks in . This vector is similar to a one-hot encoding technique; however, there is no guarantee that a pixel belongs to a single mask due to 2D segmentation inaccuracies. We visually demonstrate the formulation of in Fig. 4.
where and .
In practice, we only sample pixels from an image such that . Mathematically, we define the sampled pixel coordinates as a set .
The pairwise similarities of the sampled pixel-mask correspondence vectors can be retrieved by computing the Gram matrix of their stacked matrix . All entries of the resulting Gram matrix are then bounded to either 0 (negative pair) or 1 (positive pair) to form a mask-based correspondence matrix among sampled pixels , formally defined as
Additionally, we formulate a correspondence matrix for the sampled rendered features . Analogous to the mask-based correspondence matrix , we compute a feature-based Gram matrix, capturing the pairwise similarities between the features as follows:
where is defined as follows:
Finally, we have all the necessary ingredients to formulate the contrastive learning objective based on a hard positive and negative mining scheme. Guided by the mask-based correspondence matrix, we enforce features to be close for pixels from the same object mask and far apart for the distinct objects. Furthermore, we concentrate on hard pixel pairs. Specifically, we sample positive samples with low similarity scores and negative samples with high similarity scores. With this, we formally define the optimization objective as
Note that is an indicator function. It returns 1 when , and 0 otherwise. stands for hard positive which is defined as and refers to a hard negative which is defined as . Here, and are the positive and negative thresholds for pairwise similarity respectively. As both and are symmetric matrices following the property of the Gram matrix, we only consider entries in the upper triangular part and exclude the diagonal to avoid repeated computation. We show rendered Gaussian features learned with our contrastive scheme in Fig. 1. For demonstration, they are compressed into via PCA [33], allowing each rendered feature to be represented with RGB color.
Gaussian Feature Clustering To enable interactive applications with the learned 4D representation, we deploy the DBSCAN algorithm [12] to cluster the Gaussian features into several groups and generate the so-called Gaussian Clusters as defined in Eq. 12. provides the corresponding cluster ID for -th Gaussian feature:
Note that refers to the total number of clusters. We demonstrate them visually in Fig. 1. As can be observed, they represent distinct objects in 4D space.
Since a typical scene contains a large number of Gaussians (more than 300k), we adopt the sub-sampling scheme proposed by [6], where only 2% of the Gaussian features are used for clustering. We propagate the cluster IDs to the remaining of the Gaussians by computing the feature similarity between cluster-representative Gaussians. In the supplementary material, we demonstrate that the performance is not affected by a significant reduction in the number of Gaussians contributing to the generation of clusters.
4 Experimental Results
We evaluate the performance of SADG across several well-known datasets to validate its effectiveness. We begin by introducing the datasets in Sec. 4.1. Since there are no established benchmarks for validating the segmentation performance on novel views in dynamic scenes, we provide details on how we extend these different datasets with segmentation benchmarks (Sec. 4.2). Next, we explore the implementation details and various configurations used to train and test methods on these benchmarks (Sec. 4.3). We present segmentation results of our model, along with both quantitative and qualitative comparisons to other related works [6, 57, 25] in Sec. 4.5. Last but not least, we demonstrate capabilities of our learned semantic representation on a number of downstream tasks including scene editing and selection in our a graphical user interface (Sec. 4.6).
4.1 Datasets
NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | ||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.8351 | 0.9729 | 0.8341 | 0.9771 | 0.8864* | 0.9932* | 0.7673* | 0.9776* | 0.7979* | 0.9800* |
SAGA [6] | 0.8072 | 0.9644 | 0.7660 | 0.9579 | 0.6941 | 0.9516 | 0.7395 | 0.9747 | 0.7913 | 0.9792 |
DGD [25] | 0.7125 | 0.9551 | 0.8297 | 0.9777 | 0.7721 | 0.9851 | 0.7981 | 0.9850 | 0.7791 | 0.9728 |
SADG (Ours) | 0.8719 | 0.9826 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
NeRF-DS dataset [52] contains scenes with challenging transparent and shiny moving objects recorded by two monocular cameras. Five sequences are used for reconstruction and evaluation including as_novel_view, basin_novel_view, cup_novel_view, press_novel_view, and plate_novel_view.
HyperNeRF dataset [36] is recorded using a single monocular camera . The camera moves while capturing the dynamic scenes involving different human-object interactions. In our experiments, we use ten sequences: americano, split-cookie, oven-mitts, espresso, chickchicken, hand1-dense-v2, torchocolate, slice-banana, keyboard, and cut-lemon1.
Neu3D dataset [27] contains videos captured by a multi-view rig with to cameras each at 30 FPS. The scenes involve a person performing cooking tasks in the kitchen. In total, five scenes are used in our experiments: coffee_martini, cook_spinach, cut_roasted_beef, flame_steak, and sear_steak.
Google Immersive dataset [2] consists of videos captured by a 46-camera rig at 30 FPS. The cameras are fish-eye cameras and are distributed on the surface of a hemispherical, 92cm diameter dome. Four scenes are used in our experiments: 01_Welder, 02_Flames, 10_Alexa_1, and 11_Alexa_2.
Technicolor Light Field dataset [44] captures scenes using a camera array at 30 FPS. We use the undistorted images provided by the authors, and four scenes are used in our experiments: Birthday, Fabien, Painter, and Theater.
4.2 Segmentation Benchmarks
To date, while numerous benchmarks exist for dynamic video segmentation, such as DAVIS [38, 40, 3, 4] and VOS [51, 54, 55], there are still no well-established segmentation benchmarks for dynamic novel view synthesis.
Such benchmarks are crucial for evaluating a comprehensive understanding of scenes, both geometrically and semantically. Existing datasets, such as Replica [47], ScanNet [10], and ScanNet++ [58], include ground-truth labels for evaluating panoptic segmentation performance, but they do not cover dynamic sequences. To address this gap, we manually annotate dynamic sequences in our experiments and propose the NeRF-DS-Mask, HyperNeRF-Mask, Neu3D-Mask, Immersive-Mask, and Technicolor-Mask benchmarks. Utilizing the powerful SAM2 model [43], which provides consistent object masks across video sequences, we manually define the objects of interest for each sequence to evaluate their segmentation performance in our model. Examples of these mask objects can be seen in the supplementary.
4.3 Implementation Details
We employ SAM [20] to generate the anything-masks for Gaussian feature learning. However, the resulting masks are memory-inefficient because they are stored in a dense format () with a lot of pixels with zero values. To optimize memory usage, we save the masks as a bit array.
Given a permuted tensor version of an RGB image , SAM generates masks . We first flatten , then convert flattened ()-dimensional array into a bit-array. The final output is stored in a dictionary with the required information to re-create the original mask tensor: a number of masks , image dimensions , and the compressed bit array, i.e., . This approach significantly reduces the storage used and is particularly useful for multi-view datasets, where numerous masks from different views are generated. For instance, a raw mask generated from an image in coffee_martini occupies around 80 MB, but the dictionary representation is significantly reduced to approximately 1 MB. Due to space limits, we provide training details for different datasets in the supplementary.
4.4 Evaluation Protocol and Baselines
Given ground truth masks we use Mean Intersection over Union (mIoU) and Mean Pixel Accuracy (mAcc) to quantitatively evaluate segmentation performance of the models. We compare our model with SAGA [6], Gaussian Grouping [57] and DGD [25]. Note that both SAGA and Gaussian Grouping are designed for static scenes and do not include a module to encode temporal information. To address this limitation, we integrate our 4D reconstruction backbone into their pipeline, extending their approach to effectively handle dynamic scenes. Moreover, Gaussian Grouping needs the precomputed object IDs from DEVA [8], whose labels are inconsistent for multi-view camera setup. Thus, we train their semantic module on a monocular stream from the camera closest to the test view. Furthermore, all reported baselines utilize the same pre-trained 4D reconstruction, with semantics learned according to each pipeline. This allows us to isolate and evaluate the performance of each model based on its semantic learning capabilities. After training, we manually perform click prompts in the rendered novel views of each scene, which are trained by different approaches to select the objects of interest defined in the benchmark. Apart from the quantitative results, we also present qualitative results from various benchmarks, as certain insights cannot be fully captured through numerical evaluations alone. Since the official source code for SA4D [16] is not available, we report the comparison based on unofficial implementation in the supplementary.
4.5 Segmentation Results
As shown in Tab. 1, SADG consistently outperforms all other methods across all benchmarks in terms of average mIoU and mAcc. We also select some sequences to visually illustrate the quality of the segmentation of each model in Fig. 5. These comparisons reveal that SADG qualitatively surpasses existing models and exhibits multi-view consistent segmentation performance. Notably, SADG maintains strong generalization to test camera views in multi-view sequences such as flame_steak, Painter, and 11_Alexa_2. This robust performance can be attributed to the compact Gaussian features learned through the proposed Gaussian feature training. In contrast, DGD segments objects by comparing cosine similarity scores between features retrieved via click prompts. However, determining a suitable threshold that generalizes across various scenes is neither straightforward nor intuitive.
SADG employs clustering based on the Gaussian features to group the Gaussians into several segments. As this operation is done in 3D space, SADG achieves cross-view consistency and does not need to associate objects from different views. This advantage is evident in cut-lemon1 example in Fig. 5, where wrongly associated object IDs can cause segmentation failure for Gaussian Grouping.
4.6 Scene Editing Applications
In addition to accurate segmentation capabilities and multi-view consistent rendered masks, we demonstrate the effectiveness of our learned semantic feature field in a range of downstream tasks. Specifically, we focus on interactive editing applications and develop a graphical user interface (GUI) to grant full user control over the representation. Thanks its explicit nature and our semantic abstraction to Gaussian clusters, we achieve real-time operation capability in handling many editing tasks. We provide snapshots of the GUI in the supplementary.
Besides simple click prompts in the novel view of dynamic scenes for object selection, we also support text-based prompts to make object selection more intuitive. To enhance 3D content editing in dynamic scenes, we integrate style transfer capabilities for the selected objects. By segmenting objects directly in 3D, we facilitate scene composition and object removal through straightforward manipulation of the selected Gaussians. Due to space limit, we shall provide more qualitative results in the supplementary. In the following we describe the implementation details of editing tasks based on our semantically-aware feature field.
Text Prompt. Inspired by Gaussian Grouping [57], we employ Grounding DINO [28] to generate 2D mask based on the text prompt in the novel view. The 2D mask is reprojected into the 3D space with the rendered depth information. Finally, we associate each reprojected point with its nearest Gaussian in the 3D scene and retrieve its corresponding cluster. If the number of associated Gaussian of a given cluster is more than a threshold, the cluster would be selected. The example of text prompts on different sequences is shown in Fig. 6 (a).
Scene Composition. We also perform qualitative results on scene composition. As the segmentation is performed in 3D, we can simply put the selected Gaussians from the dynamic scene into another static or dynamic scene. The scales and coordinate transformation need to be adapted. We show some examples on scene composition in Fig. 6.
Object Style Transfer. We adopt the static Gaussian style transfer from StyleSplat [14] for dynamic scenes. Given a rendered image and a style image , we extract their feature maps (, respectively) from VGG16 [46] and optimize only the SHs of the Gaussians of the selected cluster with the nearest-neighbor feature matching (NNFM) loss [59]. Fig. 6 (c) illustrates an example of style transfer on the segmented object.
Object Removal. We can also remove object by simply deleting Gaussians belonging to the specific cluster as illustrated in Fig. 6 (d). It is worth noting that no inpainting is required due to the underlying 3D representation, where exposed parts are learned by multi-view observations.
5 Conclusion
We introduced a novel framework for dynamic scene understanding, which enables multi-view consistent segmentation without any object tracking supervision. Our SADG effectively combines dynamic 3D Gaussian Splatting 3DGS [56] and 2D SAM [20] masks in a contrastive learning objective, which lifts semantic information into 3D space and learns expressive Gaussian features based on hard pixel cases. This results in cross-view consistency when rendering segmented objects, enhancing the quality and coherence of object segmentation across different views. Evaluated on various novel-view datasets, SADG shows superior performance both quantitatively and qualitatively. We further demonstrated the effectiveness of the learned feature field on downstream editing tasks such as point and text prompts, style transfer, object removal, and scene composition. Our approach sets a strong foundation for further research into dynamic scene understanding and scene editing, especially in complex multiview scenarios.
Supplementary Material
In the supplementary, we provide further implementation details in Sec. 6 and Sec. 7. A range of ablation studies on different components of our pipeline is given in Sec. 8. We compare our SADG with the concurrent work on 4D segmentation and editing, Segment Any 4D Gaussians (SA4D), in Sec. 9. Several examples from our segmentation benchmark are given in Sec. 11. We present detailed quantitative results on all benchmarks in Sec. 10. We finally provide qualitative results of our method on segmentation (Sec. 12, Sec. 14) and editing (Sec. 13).
6 Training Implementation Details
We implement the whole pipeline in PyTorch [37]. We reuse the codebase from Deformable-3DGS [56] and extend the differential Gaussian rasterization pipeline to render the Gaussian features similar to [57]. In the following, we introduce the detailed training setup for different datasets.
6.1 NeRF-DS and HyperNeRF
For the NeRF-DS sequences, we use the left camera for training and the right camera for testing. In the HyperNeRF sequences, training frames are sampled every 4 frames, starting from the first frame, while testing frames are selected every 4 frames, starting from the third frame. This follows the interp-sequence data-loading pipeline from the original HyperNeRF paper [36], ensuring that all testing frames are unseen during training. Note that we use -downsized images for all experiments.
NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | ||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
SADG (w/o filtering) | 0.8719 | 0.9826 | 0.8368 | 0.9790 | 0.8738 | 0.9923 | 0.9154 | 0.9942 | 0.9258 | 0.9904 |
Ours (w/ filtering) | 0.8719 | 0.9826 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | ||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
SADG (K=0) | 0.8228 | 0.9706 | 0.8360 | 0.9787 | 0.8661 | 0.9920 | 0.8725 | 0.9909 | 0.9259 | 0.9908 |
Ours (K=16) | 0.8719 | 0.9826 | 0.8368 | 0.9790 | 0.8738 | 0.9923 | 0.9154 | 0.9942 | 0.9258 | 0.9904 |
SADG (K=32) | 0.8754 | 0.9828 | 0.8331 | 0.9780 | 0.8702 | 0.9922 | 0.9217 | 0.9944 | 0.9255 | 0.9904 |
Dynamic Geometry Reconstruction. We adopt the default hyper-parameters from Deformable-3DGS [56]. Initially, the Gaussians are warmed up without the deformation MLP for the first 3k iterations to better capture the positions and shape stably as suggested in [56], after which the MLP optimization begins, and the 3D Gaussian and the deformation MLP are trained jointly. The densification of Gaussians for adaptive density control introduced in 3DGS [17] is set to run until 15k iterations. The dynamic geometry reconstruction is trained for a total of 20k iterations.
Gaussian Feature Learning. After the dynamic reconstruction is completed, we freeze model’s weights and begin Gaussian feature learning for an additional 10k iterations. For each image, we randomly sample 25 masks () generated by SAM [20] and 5k pixels (), with K = 16 to get smooth Gaussian features.
6.2 Neu3D
Each video contains 300 frames, and we begin by down-scaling the images by a factor of 2, resulting in a resolution of . The center camera (camera_00) is used for testing, while the remaining cameras are for training. Initially, we run SfM [45] across all the first frames, excluding camera_00, to get the initial point cloud.
Dynamic Geometry Reconstruction. We follow the same pipeline for dynamic geometry reconstruction as described in Tab. 3, except the densification of Gaussians is set to 8k iterations.
Gaussian Feature Learning. After the dynamic reconstruction is completed, we freeze the model’s weights and begin Gaussian feature learning for an additional 10k iterations. For each image, we randomly sample 50 masks () generated by SAM [20] and 10k pixels (), with K = 16 to get the smooth Gaussian feature.
6.3 Google Immersive
Each video contains frames captured by fish-eye cameras with 300 frames. We first undistort the images and align the principal points to the center, then downscale them to a resolution of .
NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | |||||||||||
Avg. Num. Gaussians | 186455.4 | 565614.7 | 678566 | 985713 | 457677.75 | ||||||||||
Method | mIoU | mAcc | time | mIoU | mAcc | time | mIoU | mAcc | time | mIoU | mAcc | time | mIoU | mAcc | time |
Ours (2%) | 0.8719 | 0.9826 | 0.4068 | 0.8368 | 0.9790 | 2.3117 | 0.8738 | 0.9923 | 2.2506 | 0.9154 | 0.9942 | 9.6083 | 0.9258 | 0.9904 | 1.6118 |
SADG (10%) | 0.8707 | 0.9828 | 1.5782 | 0.8384 | 0.9794 | 11.8430 | 0.8792 | 0.9930 | 6.1878 | 0.9160 | 0.9939 | 71.2429 | 0.9283 | 0.9908 | 6.8211 |
SADG (25%) | 0.8762 | 0.9832 | 3.9425 | 0.8461 | 0.9809 | 43.2682 | 0.8886 | 0.9935 | 18.8976 | 0.9180 | 0.9943 | 221.7642 | 0.9297 | 0.9907 | 19.0767 |
Similar to Neu3D, the center camera (camera_0001) is used for testing, while the other cameras are used for training. We generate the initial point cloud using SfM across all the first frames, excluding the test camera, and apply the same hyper-parameters as in Neu3D in Sec. 6.2.
6.4 Technicolor Light Field
We use the undistorted images provided by the authors, aligning the principal points to the center and resizing the images to a resolution of . Similar to Neu3D and Google Immersive dataset, the center camera (camera_0000) is used for testing, while the other cameras are used for training. We generate the initial point cloud using SfM across all the first frames excluding the test camera and apply the same hyper-parameters as in Neu3D in Sec. 6.2.
7 Testing Implementation Details
After the dynamic geometry reconstruction and the Gaussian features are properly trained, we proceed to our GUI to select the object of interest and evaluate the performance of the model by mIoU and mAcc as discussed in Sec. 4.4.
Given a clicked point on the novel view, we first use the known camera intrinsics to compute the inverse intrinsic matrix . With the depth information rendered by the dynamic reconstruction of the Gaussians at time , we can reproject the 2D clicked point to 3D coordinates in the camera frame. Finally, using the known camera extrinsics (rotation matrix and translation vector ) for the world-to-camera transformation, we convert the points back into the world coordinate frame. We can then use this point coordinate to query its nearest neighbor Gaussian using KNN. The cluster ID corresponding to its nearest neighbor Gaussian is subsequently retrieved.
We provide interactive illustrations in our GUI to segment target objects as follows:
1.
Open the GUI and run DBSCAN (Click button ”Clustering”) to group Gaussians into different clusters based on the learned Gaussian features as illustrated in Fig. 7.
2.
Click the objects of interest in the novel view to segment the target objects as shown in Fig. 8.
3.
Click the button ”Render Object” to render the selected object to test views for evaluating the performance.
HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | |||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
SA4D [16] | 0.7767 | 0.9633 | 0.8832* | 0.9391* | 0.7987* | 0.9835* | 0.8271* | 0.9734* |
SADG (Ours) | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
8 Ablation Studies
8.1 Filtering
We noticed that the raw segmented object from the selected Gaussian clusters sometimes has some artifacts near the outline of the objects as illustrated in Fig. 9. Therefore, we adopt a simple yet effective filtering step to filter out those artifacts. The procedure of the filtering is as follows:
In each selected Gaussian cluster , we compute the mean Gaussian features in the selected cluster .
We compute the per-Gaussian similarity score between the Gaussian features inside the selected cluster and mean Gaussian features of the selected cluster with cosine similarity.
Delete Gaussians that have lower similarity scores than the threshold .
As shown in Tab. 2, the proposed filtering step can effectively improve the average mIoU and mAcc for each benchmark.
8.2 Smooth Gaussian Features
In Sec. 3.2 we introduced smooth Gaussian features (Eq. 6). We provide a detailed analysis of the selection of the number of nearest neighbors for the KNN algorithm in Tab. 3. Overall, when smooth Gaussian features are enabled ( and ), the performance is better than when smooth Gaussian features are disabled (). Since there is no significant improvement between and , we eventually chose for our pipeline to avoid additional overheads with the KNN sampling.
8.3 Gaussian Feature Clustering
In Sec. 3.2 we additionally introduced Gaussian Feature Clustering and the sub-sampling strategy to run the clustering in real-time with scenes that normally contain more than 300k Gaussians. We perform time analysis with different sub-sampling numbers of Gaussians (2%, 10%, and 25%) on RTX3080 and Intel i7-12700. For each scene, we compute the time for clustering using Python timeit package and take the average clustering time for five runs per scene. We report the average clustering time across all scenes in each benchmark and the corresponding mIoU and mAcc for our object of interest in Tab. 4. Overall, with different sub-sampling of Gaussians (2%, 10%, and 25%), the mIoU and mAcc stay almost unchanged. However, as increasing the number Gaussians results in a longer clustering time, we believe that 2% is the optimal sub-sampling number for clustering.
8.4 DBSCAN vs. K-Means
There are various clustering techniques and K-means [29, 32, 15, 39] is one of the most widely used. K-means partitions data into clusters, where each feature belongs to the cluster with the nearest centroid (mean of the cluster). The value must be predefined, and the algorithm iteratively adjusts the centroids to minimize the sum of squared distances between features and their corresponding cluster centers. Therefore, the performance of the algorithm heavily depends on the optimal value of .
To investigate this, we conducted an additional ablation study comparing the rendered segmentation produced by K-means with different values of and DBSCAN. As shown in Fig. 10, the choice of significantly impacts the result of rendered segmentation. Since different sequences contain varying numbers of objects, we believe that employing a method with an adaptive number of clusters, such as DBSCAN, validates our design choices with additional flexibility and no loss on the segmentation accuracy.
8.5 Time and Storage Analysis
We performed a comprehensive time analysis of SADG using an RTX 3080 GPU and an Intel i7-12700 CPU on the NeRF-DS [52] dataset. To evaluate its performance, we compared it with DGD [25], the only publicly available model addressing segmentation in dynamic novel views. The results are summarized in Tab. 6.
Most existing approaches for novel view segmentation [57, 6, 16] require precomputing SAM masks or object masks prior to training. In contrast, DGD dynamically computes DINOv2 feature maps for supervision during training. For a fair comparison, we precomputed the DINOv2 feature maps for all training images in DGD and stored them beforehand on an SSD, alongside our precomputed SAM masks.
As shown in Tab. 6, DGD [25] spends little preprocessing time for DINOv2 features (approximately 5 seconds for all training images per sequence). Although our SADG requires preprocessing time to generate SAM masks, this process only needs to be done once before training. DGD assigns a 384-dimensional semantic feature to each Gaussian. Rendering these high-dimensional features into 2D is computationally inefficient, significantly slowing down the training process. SADG uses 32-dimensional Gaussian features, which can be rendered much more efficiently. As a result, SADG achieves over 3 faster total time (preprocessing + training) than DGD while only requiring 4.5 less space in the storage.
In addition, we evaluated the processing time for click prompts in our GUI. For smaller scenes with less than 200k Gaussians such as as_novel_view, segmenting an object via a click prompt takes approximately 0.05 seconds. For larger scenes with more than 1,000k Gaussians captured by a multi-view camera setup, such as 01_Welder, the processing time per click prompt is around 0.4 seconds.
Avg. Preprocessing Time (min) | Avg. Training Time (min) | Total Time (min) | Avg. Size of Gaussians (MB) | |
DGD [25] | 0.08 | 103 | 103 | 285 |
SADG (Ours) | 16 | 16 | 32 | 60 |
9 Comparison Against Segment Any 4D Gaussians Model
We provide additional quantitative and qualitative comparisons with the concurrent work on 4D segmentation and editing, SA4D [16]. The pipeline consists of three stages: (1) dynamic 3D reconstruction based on 4D-GS model [50], (2) identity feature field learning, and (3) post-processing with outlier removal. Since the authors have not released the official code, we use the unclean version from https://github.com/guanjunwu/sa4d for training and testing the model on our benchmark. We do not provide the performance on NeRf-DS-Mask with specular and transparent moving objects in Tab. 5 since 4D-GS fails to converge on most of the sequences apart from two (basin_novel_view, press_novel_view). Similarly, we exclude two sequences from the HyperNeRF-Mask dataset, such as hand1-dense-v2 and espresso due to 4D reconstruction failure.
In quantitative assessment, we select the top 30 segmentation results to more closely reflect the test set used in the original SA4D benchmark (not released). We also uniformly sub-sample point cloud (100K points) from SfM [45]. As seen in Tab. 5, our SADG significantly outperforms SA4D on all datasets. From the qualitative results in Fig. 12 and Fig. 13, we notice that rendered masks are larger than the underlying object and include a lot of unrelated regions even after the outlier removal in the post-processing step, which negatively affects objects’ segmentation accuracy. We further demonstrate some editing comparisons in Fig. 17, Fig. 16, and Fig. 15. Due to inaccurate segmentation, composition, and removal, SA4D results contain many spurious regions, degrading the editing quality.
10 Quantitative Results per Sequence
For completeness to the average numbers per datasets (Tab. 1), we report the performance of our method individually on all benchmarks in Tab. 7, Tab. 8, Tab. 9, Tab. 10, and Tab. 11.
11 Segmentation Benchmarks
12 Additional Qualitative Results
We provide additional qualitative results in Fig. 12 and Fig. 13 to further showcase the effectiveness of SADG over other methods. While SAGA [7] suffers from degenerate Gaussians, our reconstruction backbone and effective post-processing mitigate the issue. At the same time, rendered masks from other methods, such as DGD [25] and SA4D [16], are inaccurate and introduce spurious regions. Our method demonstrates the best performance among all approaches.
13 Further Scene Editing Examples
We present further scene composite examples of SADG in Fig. 14. We also compare the quality of object removal and scene composition with SA4D [16] in Fig. 16 and Fig. 17. Overall, as discussed in Sec. 9, SADG outperforms SA4D both quantitatively and qualitatively.
as_novel_view | basin_novel_view | cup_novel_view | press_novel_view | plate_novel_view | Average | |||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.8140 | 0.9487 | 0.8083 | 0.9612 | 0.9020 | 0.9936 | 0.7803 | 0.9912 | 0.8710 | 0.9696 | 0.8351 | 0.9729 |
SAGA [6] | 0.8235 | 0.9542 | 0.6437 | 0.9140 | 0.8833 | 0.9922 | 0.8119 | 0.9918 | 0.8736 | 0.9699 | 0.8072 | 0.9644 |
DGD [25] | 0.7135 | 0.9266 | 0.6685 | 0.9291 | 0.694 | 0.9752 | 0.6457 | 0.9822 | 0.8406 | 0.9625 | 0.7125 | 0.9551 |
SADG (Ours) | 0.9162 | 0.9799 | 0.8740 | 0.9785 | 0.8864 | 0.9924 | 0.8098 | 0.9916 | 0.8733 | 0.9707 | 0.8719 | 0.9826 |
americano | chickchicken | cut-lemon1 | espresso | hand | keyboard | |||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.8980 | 0.9969 | 0.9436 | 0.9867 | 0.7153 | 0.9394 | 0.4968 | 0.9642 | 0.8865 | 0.9814 | 0.9159 | 0.9861 |
SAGA [6] | 0.7325 | 0.9873 | 0.8633 | 0.9648 | 0.5263 | 0.8398 | 0.6395 | 0.9798 | 0.9088 | 0.9859 | 0.8070 | 0.9645 |
DGD [25] | 0.7748 | 0.9912 | 0.8317 | 0.9589 | 0.8583 | 0.9718 | 0.7240 | 0.9871 | 0.8936 | 0.9838 | 0.8451 | 0.9733 |
SADG (Ours) | 0.8144 | 0.9922 | 0.9308 | 0.9835 | 0.8795 | 0.9769 | 0.7164 | 0.986 | 0.9006 | 0.9849 | 0.8952 | 0.9827 |
oven-mitts | slice-banana | split-cookie | torchocolate | Average | ||||||||
Gaussian Grouping [57] | 0.7420 | 0.9505 | 0.9095 | 0.9721 | 0.9134 | 0.9960 | 0.9196 | 0.9981 | 0.8341 | 0.9771 | ||
SAGA [6] | 0.8667 | 0.9766 | 0.7323 | 0.8960 | 0.8186 | 0.9906 | 0.7652 | 0.9933 | 0.7660 | 0.9579 | ||
DGD [25] | 0.8595 | 0.9755 | 0.8408 | 0.947 | 0.8467 | 0.9925 | 0.8223 | 0.9954 | 0.8297 | 0.9777 | ||
SADG (Ours) | 0.9295 | 0.9876 | 0.8782 | 0.9617 | 0.8581 | 0.9930 | 0.8599 | 0.9964 | 0.8663 | 0.9845 |
coffee_martini | cook_spinach | cut_roasted_beef | flame_steak | sear_steak | Average | |||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.8295 | 0.9890 | 0.8974 | 0.9941 | 0.9512 | 0.9972 | 0.8279 | 0.9900 | 0.9261 | 0.9959 | 0.8864 | 0.9932 |
SAGA [6] | 0.2201 | 0.8081 | 0.8125 | 0.9881 | 0.6982 | 0.9767 | 0.8439 | 0.9911 | 0.8959 | 0.9941 | 0.6941 | 0.9516 |
DGD [25] | 0.7875 | 0.9865 | 0.815 | 0.9883 | 0.817 | 0.9877 | 0.6771 | 0.9776 | 0.7638 | 0.9854 | 0.7721 | 0.9851 |
SADG (Ours) | 0.9120 | 0.9948 | 0.9129 | 0.9951 | 0.9103 | 0.9947 | 0.8716 | 0.9930 | 0.9044 | 0.9947 | 0.9022 | 0.9945 |
01_Welder | 02_Flames | 10_Alexa_1 | 11_Alexa_2 | Average | ||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.7920 | 0.9736 | 0.8364 | 0.9936 | 0.7391 | 0.9840 | 0.7018 | 0.9590 | 0.7673 | 0.9776 |
SAGA [6] | 0.8062 | 0.9737 | 0.8463 | 0.9940 | 0.4794 | 0.9513 | 0.8260 | 0.9797 | 0.7395 | 0.9747 |
DGD [25] | 0.873 | 0.9854 | 0.7775 | 0.9914 | 0.8107 | 0.9899 | 0.7312 | 0.9734 | 0.7981 | 0.9850 |
SADG (Ours) | 0.8975 | 0.9877 | 0.8808 | 0.9955 | 0.9367 | 0.9970 | 0.9786 | 0.9979 | 0.9234 | 0.9945 |
Birthday | Fabien | Painter | Theater | Average | ||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.9324 | 0.9908 | 0.9406 | 0.9758 | 0.8759 | 0.9888 | 0.4428 | 0.9646 | 0.7979 | 0.9800 |
SAGA [6] | 0.8541 | 0.9779 | 0.9371 | 0.9943 | 0.9670 | 0.9867 | 0.4071 | 0.9579 | 0.7913 | 0.9792 |
DGD [25] | 0.7203 | 0.9573 | 0.9177 | 0.9667 | 0.7959 | 0.9788 | 0.6826 | 0.9885 | 0.7791 | 0.9728 |
SADG (Ours) | 0.8839 | 0.9833 | 0.9734 | 0.9889 | 0.9719 | 0.9976 | 0.8941 | 0.9969 | 0.9308 | 0.9917 |
14 Visualization of the Rendered View, Rendered Segmentation, Segmented Object
We also provide the visualizations of the rendered novel view, the rendered segmentation, and the segmented objects in Fig. 18, Fig. 19, and Fig. 20. The rendered segmentation of SADG is based on Gaussian clustering, which is performed in 3D space, resulting in consistent segmentation in the dynamic scene.